### v1

# ADPro: a Test-time Adaptive Diffusion Policy for Robot Manipulation via Manifold and Initial Noise Constraints

Abstract
Diffusion policies have recently emerged as a powerful class of visuomotor controllers for robot manipulation, offering stable training and expressive multi-modal action modeling. However, existing approaches typically treat action generation as an unconstrained denoising process, ignoring valuable a priori knowledge about geometry and control structure. In this work, we propose the Adaptive Diffusion Policy (ADP), a test-time adaptation method that introduces two key inductive biases into the diffusion. First, we embed a geometric manifold constraint that aligns denoising updates with task-relevant subspaces, leveraging the fact that the relative pose between the end-effector and target scene provides a natural gradient direction, and guiding denoising along the geodesic path of the manipulation manifold. Then, to reduce unnecessary exploration and accelerate convergence, we propose an analytically guided initialization: rather than sampling from an uninformative prior, we compute a rough registration between the gripper and target scenes to propose a structured initial noisy action. ADP is compatible with pre-trained diffusion policies and requires no retraining, enabling test-time adaptation that tailors the policy to specific tasks, thereby enhancing generalization across novel tasks and environments. Experiments on RLBench, CALVIN, and real-world dataset show that ADPro, an implementation of ADP, improves success rates, generalization, and sampling efficiency, achieving up to 25% faster execution and 9% points over strong diffusion baselines.
1 Introduction
Autonomous robotic manipulation in unstructured environments requires policies that not only predict accurate actions but also generalize well across diverse tasks and scenes. Early approaches based on rule-based controllers  (Miller and Allen 2004; Li, Fu, and Pollard 2007; Dang, Weisz, and Allen 2011) or dense discriminative models  (Ten Pas et al. 2017; Liang et al. 2019; Mousavian, Eppner, and Fox 2019; Murali et al. 2020; Fang et al. 2020; Gou et al. 2021; Zhao et al. 2021; Depierre, Dellandréa, and Chen 2021; Wei et al. 2021; Ma and Huang 2023; Wei et al. 2023; Qin et al. 2023; Ryu et al. 2023) have struggled to scale, either due to lack of flexibility or the computational cost of exhaustive search. More recently, diffusion policies (DPs)  (Chi et al. 2023; Wei et al. 2025; He et al. 2025; Wang et al. 2025; Li et al. 2025; Hou et al. 2024; Ze et al. 2024; Wang et al. 2024a; Yang et al. 2024a; Ke, Gkanatsios, and Fragkiadaki 2024; Zhu et al. 2024; Yao et al. 2025; Jia et al. 2025; Cao et al. 2024) have emerged as a promising direction, enabling multi-modal action generation via iterative denoising, and achieving state-of-the-art performance on a range of manipulation benchmarks.
However, existing diffusion policies are typically trained offline and executed without test-time adaptation, using DDPM (Ho, Jain, and Abbeel 2020) or DDIM (Song, Meng, and Ermon 2021) for denoising and treating action generation as an unconstrained stochastic process (see Fig. 1.a). This overlooks key a priori knowledge and task-specific structure available at test time. For instance, the relative pose between the robot and a target object naturally suggests a direction for refinement, while coarse geometric alignment could provide better starting points. Furthermore, standard isotropic Gaussian priors inject noise uniformly across the entire action space, forcing the policy to recover from an uninformative initialization that ignores the specific characteristics of test data, resulting in inefficiencies and increased failure rates.
Figure 1: Comparison of (a) vanilla DPs and (b) our ADP.
In this paper, we propose incorporating task-specific constraints and a priori knowledge at test time, enabling a learned diffusion policy to dynamically adapt to task structure without retraining. As illustrated in Fig. 1.b, guiding the denoising process along known geometric or control-relevant manifolds (e.g., object-relative poses, contact trajectories) helps the model avoid implausible regions of the action space, resulting in more efficient inference and improved success rates. Furthermore, this structured adaptation enables the same policy to generalize across diverse tasks and environments, since it can tailor its behavior to new contexts through principled constraints rather than memorized behavior. In effect, test-time adaptation transforms a general-purpose policy into a task-specialized one on-the-fly, improving both robustness and sample efficiency.
Specifically, we introduce ADP, a test-time adaptive diffusion policy that injects structure into the denoising process through three complementary mechanisms:
  • We embed a manifold constraint guided by the relative pose between end-effector and target object, providing gradient-like guidance that steers the denoising trajectory toward task-relevant regions of the action space.
  • We refine the reverse diffusion dynamics using a Gaussian spherical prior, constraining the denoising steps to lie on a hypersphere that captures high-confidence regions of the noise distribution, reducing backtracking and improving action sampling efficiency.
  • We propose a coarse initialization strategy based on Fast Global Registration (FGR) between test-time point clouds of the gripper and target scene, ensuring the diffusion process begins from a geometrically meaningful configuration.
ADP is training-free and plug-and-play, making it readily deployable with existing pre-trained diffusion models. By implementing ADP with 3D Diffuser Actor (Ke, Gkanatsios, and Fragkiadaki 2024), we create ADPro, which improves sample efficiency, task success, and zero-shot generalization, outperforming state-of-the-art diffusion-based baselines with up to 25% fewer inference steps and 9% points improvement in success rates on long-horizon, contact-rich tasks.
2 Related Work
Diffusion Policies for Robot Control.
Diffusion policies (Chi et al. 2023; Wei et al. 2025; He et al. 2025; Wang et al. 2025; Li et al. 2025; Hou et al. 2024; Ze et al. 2024; Wang et al. 2024a; Yang et al. 2024a; Ke, Gkanatsios, and Fragkiadaki 2024; Zhu et al. 2024; Yao et al. 2025; Jia et al. 2025; Cao et al. 2024), a new type of robotic manipulation methods that combine diffusion models (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2021; Song et al. 2021; Li et al. 2023) with imitation learning, have attracted widespread attention for their ability to generate complex and realistic trajectories. These models are trained via behavior cloning to produce action sequences conditioned by visual inputs, robot states, or even language instructions. Existing approaches can be broadly categorized into two main types based on the input data type. The first one relies on 2D observations (Chi et al. 2023; He et al. 2025; Hou et al. 2024; Wang et al. 2024a; Zhu et al. 2024), typically using single or multiple RGB images. Since 2D images are heavily influenced by perspective and texture, such methods tend to be sensitive to viewpoint changes and variations in object color and texture, resulting in limited robustness. The second category leverages 3D visual information (Wei et al. 2025; Wang et al. 2025; Ze et al. 2024; Yang et al. 2024a; Ke, Gkanatsios, and Fragkiadaki 2024; Yao et al. 2025; Jia et al. 2025; Cao et al. 2024), such as point clouds or RGB-D images, offering richer environmental representations and greater generalization to variations in perspective and appearance. However, these methods face two key challenges: (1) Limited generalization to object shapes, positions, and tasks. Most approaches rely solely on gradient estimation from the diffusion model, without incorporating guidance from the target object to inform action updates. (2) Slow inference. The reverse diffusion process typically requires around 100 iterations, limiting the efficiency of action generation. Here, we address both challenges effectively by leveraging guidance from the testing data and applying spherical manifold and initial action constraints.
Robot Policy Generalization.
In robot policy learning, the achievement of generalizable policies has been a long-standing goal (Kaelbling 2020). Two main approaches are commonly used to improve controller generalization: training models on large, diverse datasets, and leveraging the broad generalization capabilities of large vision-language foundation models. For example, some researchers have expanded their training datasets by including more object instances or categories (Finn and Levine 2017; Levine et al. 2018; Depierre, Dellandréa, and Chen 2018; Young et al. 2021; Stone et al. 2023; Wang et al. 2025), novel object combinations (Dasari and Gupta 2021; Jang et al. 2022), diverse language instructions (Jiang et al. 2022; Nair et al. 2022; Wei et al. 2025), and multiple robot embodiments (Vuong et al. 2023). Given the strong representational and generalization capabilities of vision-language models (VLMs), recent works such as RT-2 (Brohan et al. 2023), RoboFlamingo (Li et al. 2024), $\pi$(Black et al. 2024a), ECoT (Zawalski et al. 2024), and Robouniview (Liu et al. 2024) have adopted pre-trained VLMs as the backbone of robotic policies to enhance robustness and generalization. Other research explores alternative perspectives, including domain adaptation (Bousmalis et al. 2018; Fang et al. 2018; Jeong et al. 2020; Ma et al. 2024a; Wu et al. 2025b, a), continuous learning (Yang et al. 2023; Wan et al. 2024; Wang et al. 2024b), and the use of prior constraints (Ma et al. 2024b; Huang et al. 2024; Yao et al. 2025). Despite recent advances, most methods focus on improving policy learning during training, while underutilizing the potential guidance from test-time data during inference. A key challenge lies in how to effectively construct guidance using observational data to improve the generalization of the policy, a problem that remains largely unexplored, and which we aim to address in this work.
3 Method
Action generation inherently involves producing non-Euclidean data, as actions reside within the $\mathrm{SE}(3)$ Lie group. Therefore, integrating a more efficient diffusion policy with a well-designed initial noisy action proposal can greatly enhance generation efficiency. As illustrated in Fig. 2, the proposed adaptive diffusion policy for manipulation action generation accelerates the process through two key improvements. First, an initial action proposal provides the diffusion process with a more reasonable and target-proximal starting action, effectively reducing the number of denoising steps required. Second, manifold constraints guide the denoising process along a more precise trajectory, preventing unnecessary backtracking (see Fig. 4). This results in diffusion trajectories with shorter geodesic distances and fewer diffusion steps, significantly improving overall efficiency.
Figure 2: The pipeline of our adaptive diffusion policy ADPro. We first propose reasonable initial noisy actions $\mathbf{a}^{M}$ with the rough registration module, then, denoise $\mathbf{a}^{M}$ under the guidance of observations and manifold constraint (see Eq. (9) for details).
3.1 Problem Statement.
Given a set of expert demonstrations containing continuous-valued action-observation pairs, denoted as $\mathcal{D}=\{(\mathbf{O}_{k},\mathbf{A}_{k})\}_{k=1}^{N}$, our objective is to fit a visuomotor policy that maps the observation space $\mathbf{O}$ to the action space $\mathbf{A}$. In this work, each observation $\mathbf{O}$ is two posed RGB-D images or point clouds. Each action $\mathbf{A}:=(\mathbf{R},\mathbf{v},w)$ is a 7D pose, defined by position $(x,y,z)$, rotation $(roll,pitch,yaw)$, and the gripper width $w$. Consequently, the pose $(\mathbf{R},\mathbf{v})$ of action lies on an $\mathrm{SE}(3)$ manifold, where $\mathbf{R}\in\mathrm{SO}(3)$ is the rotation matrix, and $\mathbf{v}\in\mathbb{R}^{3}$ is the translation vector.
Let $\mathbf{P}_{k}\in\mathbb{R}^{4\times 4}$ be the current $\mathrm{SE}(3)$ pose of the gripper in the world frame, action $\mathbf{A}_{k}$ specifies a desired pose of the gripper. The pose can be either absolute ($\mathbf{P}_{k+1}=(\mathbf{R}_{k},\mathbf{v}_{k})$, also called position control) or relative ($\mathbf{P}_{k+1}=\mathbf{R}_{k}\mathbf{P}_{k}+\mathbf{v}_{k}$, also called velocity control). To facilitate noise injection and removal using addition and subtraction, as in the standard diffusion process, we vectorize $\mathbf{A}_{k}$ into a vector $\mathbf{a}_{k}$ during diffusion and denoising. After denoising, the noise-free action vectors are converted back into $\mathbf{R}$, $\mathbf{v}$, and $w$ for manipulation.
For trajectory generation, $\mathbf{O}=\{\mathbf{O}_{k-(m-1)},\dots,$ $\mathbf{O}_{k-1},\mathbf{O}_{k}\}$, $\mathbf{a}=\{\mathbf{a}_{k},\mathbf{a}_{k+1},\dots,\mathbf{a}_{k+(n-1)}\}$ where $m$ is the number of history steps observed and $n$ is the number of future action steps. Given that diffusion policies involve both multiple-action trajectories across different positions and a single denoising trajectory for each action at a given position, we introduce a clear notation system to avoid ambiguity. We use a subscript to denote individual action $\mathbf{a}_{k}$ within action trajectories and a superscript to represent noisy action trajectory $\mathbf{a}^{t}$ during denoising process for a single action.
3.2 Preliminaries
Diffusion Policy.
[…]
$\mathbf{a}^{t-1}_{k}=\alpha(\mathbf{a}^{t}_{k}-\gamma\varepsilon_{\theta}(\mathbf{O}_{k},\mathbf{a}^{t}_{k},t)+\epsilon),$ (1)
where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. $\alpha,\gamma,\sigma$ are functions of the denoising step $t$ (also known as the noise schedule). The action $\mathbf{a}^{0}$ is expected to be the corresponding action from the expert policy $\pi:\mathbf{O}\mapsto\mathbf{a}$.
Training-Free Guided Diffusion Models.
Classifier guidance (Dhariwal and Nichol 2021) is the first work that utilizes the pre-trained diffusion model for conditional image generation in a training-free fashion. Specifically, considering the Bayes rule $p(x|y)=p(y|x)p(x)/p(y)$, it introduces the given condition with an additional likelihood term $p(x_{t}|y)$:
$\nabla_{x_{t}}\log p(x_{t}|y)=\nabla_{x_{t}}\log p(x_{t})+\nabla_{x_{t}}\log p(y|x_{t}),$ (2)
where $y$ is the class label, $x_{t}\sim\mathcal{N}(\sqrt{\bar{\alpha}_{t}}x_{0},\sqrt{1-\bar{\alpha}_{t}}I)$, $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$. Without training the time-dependent classifier to estimate $p(y|x_{t})$, these training-free guided diffusion models (Dhariwal and Nichol 2021; Chung et al. 2023; Yang et al. 2024b) just need a pre-trained diffusion prior $\varepsilon_{\theta}$ and a differentiable loss $L(x_{0},y)$ which is defined on the support of $x_{0}$. They use Tweedie’s formula to calculate $\nabla_{x_{t}}\log p(y|x_{t})$ by estimating $\hat{x}_{0}$ based on $x_{t}$:
$\begin{split}\hat{x}_{0}(x_{t})&\approx\mathbb{E}[x_{0}|x_{t}]=(x_{t}+b_{t}^{2}\nabla_{x_{t}}\log p(x_{t}))/a_{t},\\ \nabla_{x_{t}}\log p(y|x_{t})&\approx\nabla_{x_{t}}\log p(y|\hat{x}_{0}(x_{t}))=\eta\nabla_{x_{t}}L(\hat{x}_{0}(x_{t}),y).\end{split}$ (3)
Then they use the estimated likelihood of $L(x_{0},y)$ for additional correction step:
$x_{t-1}=\underbrace{DDIM(x_{t},\varepsilon_{\theta}(x_{t},t),t)}_{\text{sampling step}}-\underbrace{\eta\nabla_{x_{t}}L(\hat{x}_{0}(x_{t}),y)}_{\text{correction step}}.$ (4)
3.3 Adaptive Diffusion Policy with Manifold Constraint
Guided diffusion models have achieved significant success in generating Euclidean data, such as images. Given the limitations of diffusion policies in generalizing across different tasks, we believe that training-free guided diffusion policies offer an effective solution. However, existing guided diffusion models primarily focus on Euclidean data generation, and in the domain of robotic manipulation, constructing effective guidance remains a challenge.
In robotic manipulation, the vanilla DPs takes $\varepsilon_{\theta}(\mathbf{O}_{k},\mathbf{a}_{k}^{t},t)$ as the denoising direction, serving as the primary update direction for action refinement. However, it does not explicitly leverage the target object to optimize the action more explicitly and effectively. Actually, the relative position between the target object and the end-effector naturally provides the gradient direction for refining the action, directly guiding the end-effector toward objects. Therefore, we integrate the observation guidance $\nabla_{\mathbf{a}}L(\cdot,\cdot)$ into diffusion process and obtain
$\tilde{\mathbf{a}}_{k}^{t-1}$ $=DDPM(\mathbf{a}^{t}_{k},\varepsilon_{\theta}(\mathbf{O}_{k},\mathbf{a}^{t}_{k},t),t)\,$, (5)
$\mathbf{a}_{k}^{t-1}$ $=\tilde{\mathbf{a}}_{k}^{t-1}-\eta\nabla_{\mathbf{a}}L(\tilde{\mathbf{A}}_{k}^{t-1}\mathbf{O}_{k,0},\mathbf{O}_{k,1})\,.$ (6)
Here, $\mathbf{O}_{k,0}$ and $\mathbf{O}_{k,1}$ represent the gripper and scene point clouds extracted from $\mathbf{O}_{k}$, respectively. $\tilde{\mathbf{A}}_{k}^{t-1}$ denotes the matrix form of $\tilde{\mathbf{a}}_{k}^{t-1}$, while $L(\cdot,\cdot)$ is Chamfer distance.
However, since $\tilde{\mathbf{A}}_{k}^{t-1}$ is a noisy action rather than the final action executed by the robot, the guidance provided by Eq. (6) is not sufficiently accurate. Inspired by DDPM (Ho, Jain, and Abbeel 2020), we obtain the noise-free action $\hat{\mathbf{a}}_{k}^{0}$ from $\tilde{\mathbf{a}}_{k}^{t-1}$ using the diffusion function, and reformulate Eq. (6) as
$\hat{\mathbf{a}}_{k}^{0}=(\tilde{\mathbf{a}}_{k}^{t-1}-\sqrt{1-\bar{\alpha}_{t}}\varepsilon_{\theta}(\mathbf{O}_{k},\mathbf{a}^{t}_{k},t))/\sqrt{\bar{\alpha}_{t}}\,$, (7)
$\mathbf{a}_{k}^{t-1}=\tilde{\mathbf{a}}_{k}^{t-1}-\eta\nabla_{\mathbf{a}}L(\tilde{\mathbf{A}}_{k}^{0}\mathbf{O}_{k,0},\mathbf{O}_{k,1})\,.$ (8)
With the observation guidance $\nabla_{\mathbf{a}}L(\cdot,\cdot)$, Eq. (8) constrains the update direction of the inverse diffusion process and shortens the diffusion trajectory. Moreover, by incorporating observation point clouds from the test phase into the guidance, Eq. (8) effectively enhances the generalization of DPs. However, it implicitly relies on a strong assumption—the linear manifold, which may lead to manifold deviation and unnecessary backtracking in the trajectory.
To minimize unnecessary backtracking in the diffusion process (illustrated as Diffuser in Fig. 4), we introduce a manifold constraint into the reverse diffusion policy, ensuring that the denoising updates align more closely with the manifold’s geodesic path. Specifically, we adopt the Gaussian spherical prior (Yang et al. 2024b) to refine the reverse diffusion step size. The Gaussian spherical prior constrains the noise sampling space to a uniform distribution on the hypersphere of radius $\sqrt{d}\sigma$, keeping updated step size within high-confidence intervals. Here, $d$ is the product of the action dimension, the batch size, and the number of actions $n$ to be predicted. As remarked in (Yang et al. 2024b), when $d$ is sufficiently large, a $d$-dimensional isotropy Gaussian distribution $\mathcal{N}(0,\sigma^{2}I)$ is close to a uniform distribution on the hypersphere of radius $\sqrt{d}\sigma$. Then, the refined training-free adaptive diffusion policy is formulated as:
[…]
3.4 Initial Noise Constraints by Action Proposal
Given the end-effector $\mathbf{O}_{k,0}$ and a target scene $\mathbf{O}_{k,1}$, the search space for initial actions should be restricted to a specific local region rather than spanning the entire noise space, as is common in existing diffusion policies. Relying on a predefined noise distribution for sampling across the full noise space not only increases the complexity of subsequent denoising but also lowers the manipulation success rate. A more effective way is to propose a reasonable initial action by computing a coarse registration, which simplifies denoising and reduces the number of diffusion steps needed.
Specifically, we employ Fast Global Registration (FGR) algorithm (Zhou, Park, and Koltun 2016) to achieve coarse registration between the observed data $\mathbf{O}_{k,0}$ and $\mathbf{O}_{k,1}$. The resulting output serves as the proposed initial noisy action at step $M$ ($M\leq T$). Compared to the classical Iterative Closest Point (ICP) (Besl and McKay 1992), FGR offers improved efficiency and robustness, particularly when dealing with noisy or large-scale point clouds.
$\mathbf{a}_{k}^{M}=FGR(\mathbf{O}_{k,0},\mathbf{O}_{k,1})$ (10)
Figure 3: Trajectory comparison between the vanilla diffusion policy and our ADPro. We illustrate two distinct tasks: ‘put groceries in cupboard’ and ‘insert peg’. For each task, we visualize the initial state followed by successive observations captured from the wrist-mounted camera after each action. ADPro completes the first task with only three actions, demonstrating both efficiency and accuracy.
3.5 Implementation and Error Analysis of ADP
For the robotic manipulation, ADP can generate control actions and be seamlessly integrated with other DPs in a plug-and-play fashion, without requiring model retraining. Specifically, we first generate noisy initial actions $\mathbf{a}_{k}^{T}$ using Eq. (10), and then iteratively denoise them with Eq. (5), Eq. (7) and Eq. (9) until the final noiseless action $\mathbf{a}_{k}^{0}$ is obtained.
Table 1: The mean success rate on RLBench across three random seeds. ADPro outperforms all prior arts on most tasks by a large margin. Bold indicates the best; underlining shows the suboptimal.
Avg. open slide sweep to put in turn put in place stack sort insert drag place close meat off stack push screw put in Success $\uparrow$ drawer block dustpan safe tap drawer wine cups shape peg stick cups jar grill blocks buttons bulb cupboard C2F-ARM 20.1 20 16 0 12 68 4 8 0 8 4 24 24 0 20 0 72 8 0 PerAct 49.4 88.0 74.0 52.0 86.0 88.0 51.2 44.8 2.4 16.8 5.6 89.6 2.4 55.2 70.4 92.8 0 17.6 28.0 HiveFormer 45 52 64 28 76 80 68 80 0 8 0 76 0 52 100 8 84 8 32 PolarNet 46 84 56 52 84 80 32 40 8 12 5 92 0 36 100 4 96 44 12 RVT 62.9 71.2 81.6 72.0 91.2 93.6 88.0 91.6 26.4 36.0 11.2 99.2 4.0 52.0 88.0 28.8 100 48.0 49.6 Act3D 63.2 78.4 96.0 86.4 95.2 94.4 91.2 59.2 9.6 29.6 24.0 80.8 3.2 96.8 95.2 4.0 93.6 32.8 67.2 Diffuser 81.3 89.6 97.6 84.0 97.6 99.2 95.0 93.6 47.2 44.0 65.6 100.0 24.0 96.0 96.8 66.8 98.4 82.4 85.6 VPDD 56.2 - 70.7 40.0 70.7 88.7 24.0 60.0 56.0 - 73.7 66.7 30.7 37.3 73.3 56.0 58.7 61.3 30.67 Ours 83.9 93.1 100.0 97.0 98.5 99.5 95.2 94.0 56.5 50.0 74.5 100.0 25.4 96.2 95.6 67.3 98.6 83.1 85.9
Figure 4: Comparison on components ($x$, $z$, $r_{1}$, $r_{2}$) of the full action for task ‘sweep to dustpan’. Full action comparisons are available in the supplementary. The horizontal and vertical axes represent action steps and parameter values, respectively. Our ADPro effectively mitigates coordinate and angle backtracking behaviors in the vanilla diffusion policy.
Figure 5: Diffusion-step evolution for the first four actions in task ‘sweep to dustpan’. The MSE of Diffuser exhibits large fluctuations, indicating significant backtracking behavior in its trajectory.
Considering that 3D robot policies have shown superior generalization over 2D policies across varying camera viewpoints and are more effective in handling novel viewpoints during testing (Ke, Gkanatsios, and Fragkiadaki 2024; Ze et al. 2024; Jia et al. 2025), we implement ADP with 3D Diffuser Actor to create ADPro. ADPro’s conditional noise predictor $\varepsilon_{\theta}$ , leverages a pretrained 3D Transformer-based network. To further enhance the model’s generalization across diverse tasks, we incorporate task instructions as textual prompts encoded by the text encoder of the vision-language model CLIP (Radford et al. 2021). The resulting text features are then fused with visual features through a cross-attention mechanism, enabling a more effective integration of semantic and visual information.
Error Analysis.
The robustness of the proposed ADP depends on both the FGR algorithm and the pre-trained diffusion model. When $\mathbf{a}_{k}^{M}$ in Eq. 10 satisfies $\mathbf{a}_{k}^{M}-\mathbf{a}_{k}^{0}||<\delta_{1}$, the noisy action $\hat{\mathbf{a}}_{k}^{M}$ of the pretrained $\varepsilon_{\theta}$ at steps $M$ satisfies $\hat{\mathbf{a}}_{k}^{M}-\mathbf{a}_{k}^{0}||<\delta_{2}$, then, $\mathbf{a}_{k}^{M}-\hat{\mathbf{a}}_{k}^{M}||<\delta_{1}+\delta_{2}$. Furthermore, $\delta_{1}$ is controllable and can be reduced by increasing the number of iterations of the FGR algorithm, which in turn lowers $\delta_{1}$ and reduces the error upper bound. This ensures the robustness of the proposed ADP.
4 Experiments
We designed experiments to answer the following questions: i) Does the proposed diffusion policy improve manipulation performance? ii) How does ADPro perform in terms of generalization and efficiency? iii) Can ADPro be applied to real-world robotic manipulation tasks? We first demonstrate the performance improvements of our proposed ADPro on RLBench (James et al. 2020) and CALVIN (Mees et al. 2022). Then, in Sections 4.3 and 4.4, we validate its enhancements in generalization and efficiency on CALVIN and a real-world dataset. Finally, we conduct ablation studies in Section 4.5 to verify the effectiveness of each core module.
4.1 Baselines
We evaluate the proposed methods with thirteen baselines in four datasets, including C2F-ARM (James et al. 2022), PerAct (Shridhar, Manuelli, and Fox 2023), HiveFormer (Guhur et al. 2023), PolarNet (Chen et al. 2023), RVT (Goyal et al. 2023), Act3D (Gervet et al. 2023), HULC (Mees, Hermann, and Burgard 2022), RoboFlamingo (Li et al. 2024), SuSIE (Black et al. 2024b), GR-1 (Wu et al. 2024), and three DPs: 3D Diffusion Policy (DP3) (Ze et al. 2024), 3D Diffuser Actor (Diffuser) (Ke, Gkanatsios, and Fragkiadaki 2024), VPDD (He et al. 2025). For VPDD, we use the results reported in their official paper. Since no publicly available pretrained model of DP3 for CALVIN exists, we train it on CALVIN to ensure a fair comparison. For the remaining baselines, we adopt the results from 3D Diffuser Actor. Firstly, we leverage benchmarks RLBench and CALVIN to present the effectiveness of the proposed ADP on task success rates and inference speed. Meanwhile, we demonstrate the generalization of our policy in real-world scenes with Franka Panda Robot. We also compare against ablative versions that do not consider manifold or initial noise constraints.
4.2 Performance on RLBench
RLBench is a simulation environment where a Franka Panda Robot is used to manipulate the scene with BiRRT (Kuffner and LaValle 2000) as its motion planner. On RLBench, all methods need to predict the next end-effector keypose. We evaluate ADPro on 18 tasks as  (Ke, Gkanatsios, and Fragkiadaki 2024), each with 2-60 variations. We evaluate policies by task completion success rate, the ratio of execution trajectories that achieve the goal conditions specified in the language instructions (Gervet et al. 2023; Shridhar, Manuelli, and Fox 2023). Tab. 1 presents the evaluation metrics. The baseline performances are reported from Diffuser, and our results are averaged over three random seeds.
As shown in Table 1, our method outperforms the vanilla diffusion policy (Diffuser), with particularly notable improvements on the tasks ‘sweep to dustpan’, ‘sort shape’, and ‘insert peg’. Additionally, we show the robot arm’s Cartesian pose and gripper-gap evolution during a ‘sweep to dustpan’ episode in Fig. 4. From the plots of $r_{2}$ and $r_{3}$, it can be seen that our approach yields more consistent values, and it completes the task using only four actions—one fewer than the five used in the ground-truth demonstration. This reduction in action count stems from our guided diffusion mechanism, which accelerates the grasping process. In Fig. 5, we examine the inverse diffusion process for each action prediction. To enhance readability, we compute the Mean Squared Error (MSE) between the predicted actions and the ground truth, where values closer to zero indicate better predictions. The results demonstrate that incorporating our guidance during generation substantially reduces unnecessary backtracking and improves stability, yielding a much smoother generation trajectory. Notably, for ‘Action 2’ and ‘Action 4’, due to alignment, the initial position of the diffusion process is closer to the ground truth, which facilitates faster convergence toward an accurate prediction and reduces the number of diffusion steps required. In the left panel of Fig. 6, we vary the number of diffusion steps used for generation and evaluate the average success rate on RLBench. We find that our method significantly outperforms the baseline, particularly when using a small number of diffusion steps. Notably, our method achieves a $72\%$ success rate with only 20 steps, whereas the baseline requires at least 50 steps to reach comparable performance. Moreover, our method consistently improves upon the baseline across all diffusion step settings.
Table 2: Zero-shot long-horizon evaluation on CALVIN.
Method Task completed in a row Avg. Avg. Time(s)
1 2 3 4 5 Len $\uparrow$ / NDS $\downarrow$
HULC 41.8 16.5 5.7 1.9 1.1 0.67
RoboFlamingo 82.4 61.9 46.6 33.1 23.5 2.48
SuSIE 87.0 69.0 49.0 38.0 26.0 2.69
GR-1 85.4 71.2 59.6 49.7 40.1 3.06
DP3 53.9 44.7 38.0 34.3 29.0 2.00 36.1 / 25
DP3+ADP 78.4 63.7 54.9 39.8 30.2 2.67 27.3 / 10
Diffuser 93.8 80.3 66.2 53.3 41.2 3.35 48.3 / 25
Ours 94.7 83.0 73.6 61.4 51.1 3.64 36.2 / 10
4.3 Generalization on CALVIN
The CALVIN benchmark is built on the PyBullet (Coumans and Bai 2016) simulator and features a Franka Panda robot arm interacting with its scene. It includes 34 tasks across four environments (A, B, C, and D), each equipped with a desk, sliding door, drawer, LED button, lightbulb switch, and three colored blocks. The environments vary in desk texture and object placement. CALVIN provides 24 hours of unstructured play data, with 35% annotated using natural language.
We evaluate models under the zero-shot generalization setup: training on environments A, B, and C, and testing on D. Since CALVIN lacks a motion planner, models must directly predict robot pose trajectories. For a fair comparison, we follow the same evaluation protocol as Diffuser. Each method predicts at most 60 actions per sequence and is evaluated across three random seeds. We report the success rate and the average number of sequential tasks completed. Additionally, we compare DP3 and Diffuser in terms of task completion time (in seconds) and the number of diffusion steps (NDS) required, with results shown in the final column. Other methods are excluded from this comparison, as they do not use diffusion-based policies.
As shown in Tab. 2, our method (last row) consistently outperforms existing approaches on the zero-shot long-horizon prediction task, achieving a roughly 10-point improvement in the success rate of the 5th key pose and producing trajectories with the highest average length. Furthermore, pairwise comparisons between DP3 and DP3+ADP (which integrates our proposed ADP into DP3), and between Diffuser and our method (Diffuser with ADP) highlight the effectiveness of ADP as a plug-and-play module. ADP significantly improves the generalization performance of DPs and boosts generation efficiency, reducing time costs by approximately 25%.
Figure 6: Evaluating the influence of diffusion steps on two benchmarks: RLBench and Acronym.
Table 3: Evaluation of generalizability of ADPro on RealWP for five continuous key pose prediction, with and without fine-tuning.
Methods with fine-tuning Avg. Len without fine-tuning Avg. Len Avg. Time(s)
1 2 3 4 5 1 2 3 4 5
Diffuser 66.0 49.5 36.6 26.8 18.5 1.97 38.8 26.5 16.3 8.2 2.1 0.92 0.95
Ours 86.5 67.5 52.1 39.0 27.7 2.73 53.1 40.8 32.7 24.5 16.3 1.18 0.85
Table 4: Comparison of grasp success and inference time across 4 Acronym object categories.
Bottle Laptop Book ToyFigure
Avg. Success Avg. Time Avg. Success Avg. Time Avg. Success Avg. Time Avg. Success Avg. Time
SEDF(S=20) 48.0 0.019 37.1 0.028 61.2 0.020 38.6 0.024
SEDF(S=120) 51.4 0.133 41.9 0.116 66.8 0.113 44.0 0.110
SEDF+ADP(S=20) 52.7 0.025 43.0 0.032 73.1 0.027 44.0 0.029
4.4 Generalization on Real-world Data
Real_World_Play (RealWP) (Mees, Borja-Diaz, and Burgard 2023) contains 9 hours of unstructured real-world data collected by experts using Franka Panda robots. The dataset covers over 25 distinct manipulation skills, which are highly diverse and challenging due to their unstructured and sometimes suboptimal nature. A total of 3,605 episodes are annotated with language descriptions of the robot’s behavior. Each episode includes RGB-D images from a static and a gripper-mounted camera, proprioceptive data, and a 7-DoF action. Since RealWP lacks a motion planner, we predict gripper pose trajectories, similar to CALVIN. To do this, we extract frames with significant motion changes as key poses using the algorithm from 3D Diffuser Actor. This process yields a trajectory of 6, 7, or 8 steps for each episode.
To evaluate the generalization ability of ADP, we directly applied the model trained on CALVIN to predict five key poses sequentially. The accuracy of consecutive five-step trajectory, shown on the right side of Tab. 3, demonstrates that our method enhances generalization across tasks and environments, even without fine-tuning. We further fine-tuned the model using 80% of the 3,605 episodes and evaluated it on the remaining 20%. The results, presented on the left Tab. 3, show that after fine-tuning, our ADPro substantially boosts the policy’s one-step success rate by 20 percentage points.
4.5 Ablations
The proposed adaptive policy can be seamlessly integrated into existing diffusion policies in a plug-and-play manner. To assess its effectiveness in improving success rate and efficiency of manipulation, we incorporated it into three baseline models: DP3, SE(3)-DiffusionFields (SEDF) (Urain et al. 2023), and Diffuser. We then conducted a quantitative comparison of their results, as shown in Tab. 2 and 4.
For SEDF, we evaluate the pre-trained model on the Acronym dataset (Eppner, Mousavian, and Fox 2021), which provides successful 6-DoF grasps for diverse ShapeNet (Chang et al. 2015) objects. Grasp success rates are evaluated in simulation using Isaac Gym (Makoviychuk et al. 2021), following the same evaluation protocol of (Urain et al. 2023). We evaluate the model on four representative object categories: Bottle, Laptop, Book, and ToyFigure, each with the first 50 instances selected from the dataset. For each category, we report the average grasp success rate and the average generation time per grasp (in seconds), under both low-step (S=20) and full-step (S=120) diffusion settings. The results are summarized in Tab. 4. Our method (SEDF+ADP) achieves either comparable or superior success rates compared to both the baseline SEDF (S=20) and the high-step SEDF (S=120), while achieving up to 5× faster grasp generation compared to SEDF (S=120). In addition, we evaluate the multi-object model on a diverse set of 61 objects, each from a distinct category, by executing 200 generated grasps per object. As shown in Fig. 6, integrating ADP consistently improves performance across all diffusion step settings. On average, our method achieves a 2.7% increase in success rate compared to SEDF with the same steps.
Furthermore, to evaluate the effectiveness of the major component of our ADPro, we consider the following ablative versions of our model: i) without (w/o) initial noise constraint (INC), ii) w/o spherical Gaussian constraint (SGC), iii) w/o observation guidance (OG). Ablation studies on the RLBench and CALVIN benchmarks show that all three modules contribute effectively to model performance. Notably, removing the OG module leads to a substantial drop in performance and increased time costs, as shown in Tab. 5.
Table 5: Ablation study. Our model significantly outperforms its counterparts that do not use the initial noise constraint, the Gaussian manifold constraint, or the observation guidance.
Versions RLBench CALVIN
Avg. Succ. Avg. Time(s) Avg. Len Avg. Time(s)
w/o INC 82.4 421.6 3.60 35.2
w/o SGC 82.1 418.4 3.57 35.8
w/o OG 81.6 488.7 3.38 49.3
ADPro 83.9 435.8 3.64 36.2
5 Conclusion
In this paper, we present ADPro, a novel and adaptive diffusion policy that leverages observation guidance from observed data and constrains the diffusion process to improve generalizability in robotic manipulation. Unlike traditional methods, ADPro can generate transferable actions for unseen tasks without retraining. By imposing a spherical manifold constraint, the diffusion trajectories exhibit reduced backtracking and improved efficiency. Additionally, by initializing actions using the FGR algorithm rather than random noise, ADPro benefits from more informed starting points. Extensive experiments demonstrate that our training-free approach outperforms existing methods in both generalization and efficiency, confirming the potential of our approach.
The limitation of our proposed method is its reliance on both the end-effector and the scene point clouds as input, which may constrain its applicable scenarios. However, such data is readily available in some datasets and can also be feasibly acquired on real robotic platforms. In future work, we will explore more flexible forms of guidance to reduce dependency on specific input modalities.